Mixing investments. Helen invested $40,000 and received a total of $ 2300 in interest after one year. If part of the money returned 5% and the remainder 8%, then how much did she invest at each rate?
Answer by nerdybill(7384)

Mixing investments. Helen invested $40,000 and received a total of $ 2300 in interest after one year. If part of the money returned 5% and the remainder 8%, then how much did she invest at each rate?
.
Let x = amount invested at 5%
then
40000-x = amount invested at 8%
.
Because we were given the total interest earned:
2300 = .05x + .08(40000-x)
2300 = .05x + 3200 - .08x
2300 = 3200 - .03x
2300 + .03x = 3200
.03x = 900
x = $30,000 (invested at 5%)
.
Amount invested at 8%:
40000-x = 40000-30000 = $10,000 (invested at 8%)